HanAkoSolo (jamierox4ev3r):

Find the exact value of:

HanAkoSolo (jamierox4ev3r):

\(\Large\cos (-\frac{5\pi}{4})\)

HanAkoSolo (jamierox4ev3r):

the negative sign is throwing me off

OpenStudy (zzr0ck3r):

\(\cos(-x)=\cos(x)\)

HanAkoSolo (jamierox4ev3r):

wait. so is it basically just saying \(\Large\cos (\frac{5\pi}{4})\) o-o

OpenStudy (zzr0ck3r):

correct

OpenStudy (matlee):

\[-\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (zzr0ck3r):

@matlee we all know how to use a calculator.

OpenStudy (matlee):

your smart , i hope you do

OpenStudy (zzr0ck3r):

read the rules, if you can

OpenStudy (bradely):

cos(-5pi/4) = cos(5pi/4) =cos(2pi -3pi/4) =cos(3pi/4) =cos(pi-pi/4) =-cospi/4 =-1/sqrt(2) Source: mathskey.com

HanAkoSolo (jamierox4ev3r):

@matlee and @bradely , please don't provide direct answers, I don't care what the source is. So from what i see, \(\Large\cos (-\frac{5\pi}{4}) = -\frac{\sqrt{2}}{2}\)

HanAkoSolo (jamierox4ev3r):

So what was the purpose of putting a negative sign in there? Just to throw of poor unsuspecting students? Or is there actually something I need to do about the said negative sign?

HanAkoSolo (jamierox4ev3r):

@Nnesha @zzr0ck3r

HanAkoSolo (jamierox4ev3r):

and @satellite73 thank you for the trig cheat sheet, lot of helpful things on there

Nnesha (nnesha):

cos is even function cos(-x)= cos(x) but careful sin and tan sin(-x)=-sin(X) odd tan(-x)=-tan(x)odd

Nnesha (nnesha):

about*

Nnesha (nnesha):

so if cos (-x)= cos(x) then sec(-x)= ?

HanAkoSolo (jamierox4ev3r):

then sec(-x) would be sec(x), since sec is just the inverse function of cos

Nnesha (nnesha):

so -cos(-x)= ?

HanAkoSolo (jamierox4ev3r):

-cos(-x)= -cos(x)

Nnesha (nnesha):

yep! there is an identity to prove cos(-x)=cos(x) you'll learn n calc one i guess

HanAkoSolo (jamierox4ev3r):

funnnn

Nnesha (nnesha):

ye!

HanAkoSolo (jamierox4ev3r):

XD Thank you

Nnesha (nnesha):

np :=)

OpenStudy (zzr0ck3r):

cos gives you the \(x\) value on the angle,  The bottom angle is the negative version of the top angle, but they both give the same x values.

OpenStudy (zzr0ck3r):

I hope that makes sense.